How do I solve the square root of x+6 =-9?
Found 2 solutions by lynnlo, Alan3354:
Answer by lynnlo(4176)

in your problem if that is √x+6=(-)minus 9 ,there is no solution
if it is √x+6=9
the answer is ,x=9

Answer by Alan3354(69443)

How do I solve the square root of x+6 =-9?
------
Square both sides.